In her first year as a dog breeder, her dogs produced 2 puppies. In her second year her dogs produced three times as many puppies as the previous year. In her third year her dogs produced five times as many puppies as the first year. How many puppies have been produced in three years?

Answer by rfer(16322)

2+3*2+5*2=
2+6+10=18